# Audio–Visual Speech Enhancement: Architectural Design and Deployment Strategies

### v1

@@ Abstract
This paper introduce a new AI-based Audio-Visual Speech Enhancement (AVSE) and presents a comparative performance analysis of deployment architectures. The proposed AVSE system integrates convolutional neural networks (CNNs) for spectral feature extraction and long short-term memory (LSTM) networks for temporal modeling, enabling robust speech enhancement through multimodal fusion of audio and visual cues. We investigate multiple deployment scenarios, including cloud-based, edge-assisted, and standalone device implementations, evaluating their performance in terms of speech quality improvement, latency, and computational overhead. Real-world experiments are conducted across heterogeneous network conditions (Ethernet, Wi-Fi, 4G, and 5G) to assess the trade-offs between processing delay, communication latency, and perceptual speech quality. Results show that while cloud deployment achieves superior enhancement quality, edge-assisted architectures offer the best compromise between latency and intelligibility, meeting real-time requirements under 5G and Wi-Fi 6 conditions. The findings provide practical guidelines for selecting and optimizing AVSE deployment architectures in diverse application domains, from assistive hearing devices to telepresence and industrial communications.
@@
Audio–Visual Speech Enhancement, Deep Learning, Deployment Architecture, Comparative Performance, COG-MHEAR, Edge Computing, Cloud Processing.
@@ I INTRODUCTION
Audio-Visual Speech Enhancement (AVSE) leverages the complementary nature of auditory and visual cues to improve speech perception in adverse acoustic conditions. In humans, this multimodal integration is an innate cognitive process, enabling effective communication in the presence of background noise, competing talkers, or reverberant environments. Neuroscientific studies have established the value of visual input in selective auditory attention. Golumbic et al. [13] demonstrated that visual access to a talker’s face significantly improves the auditory cortex’s ability to segregate and track the target speaker in multi-speaker environments, a phenomenon often referred to as the “Cocktail Party” effect. Likewise, McGurk and MacDonald [3] revealed that incongruent lip movements can alter auditory perception entirely, demonstrating that speech understanding is not purely an auditory process but is strongly influenced by visual context. Further, Partan and Marler [9] emphasized that multimodal communication signals, common in group-living species, improve clarity, robustness, and interpretation.
Motivated by these findings, AVSE systems employ visual information—such as lip motion, facial dynamics, and articulatory patterns—to guide audio enhancement. The integration of these modalities allows for more accurate speaker separation, noise suppression, and reverberation reduction. Modern deep learning techniques have accelerated this field by enabling hierarchical feature extraction and cross-modal fusion at multiple abstraction levels. Convolutional Neural Networks (CNNs) are typically employed to extract discriminative visual features, while spectrogram-based CNNs or waveform-level encoders capture audio representations. Recurrent architectures, such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs), are frequently used to model temporal dependencies, which are critical for capturing speech dynamics.
The applications of AVSE extend beyond traditional hearing aids. In military communications, AVSE can enhance intelligibility in complex acoustic environments such as the battlefield. In consumer electronics, it improves voice-command recognition for smart assistants and in-car systems, even in noisy driving conditions [27]. In education and telepresence, AVSE can ensure intelligible communication in virtual classrooms or video conferencing, while in industrial and public-service robotics, it enables reliable interaction despite background noise. The healthcare sector can also benefit, as AVSE improves telemedicine consultations for patients with speech or hearing impairments [26].
While algorithmic design is central to AVSE performance, deployment architecture has a decisive role in real-world effectiveness. Processing location—whether cloud-based, edge-assisted, or entirely on-device—affects latency, speech quality, and system scalability. Communication bandwidth, device computational capability, and network stability determine whether a deployment can meet the real-time requirements for interactive speech enhancement. In particular, deployment trade-offs become critical when serving resource-constrained devices such as IoT endpoints, low-cost smart glasses, or wearable hearing-assist devices.
In this paper, we present a complete AVSE system implemented within the COG-MHEAR framework. The system uses a CNN–LSTM-based fusion architecture, and we evaluate its performance across multiple deployment configurations: server-side cloud processing, edge-assisted processing, and local execution. The contributions of this work are threefold:
  1. 1.
    We propose an AVSE architecture that combines convolutional and recurrent feature extraction with intermediate fusion, enabling robust speech enhancement in challenging conditions.
  2. 2.
    We design and implement a flexible deployment framework capable of supporting heterogeneous device capabilities, from lightweight IoT hardware to high-performance computing platforms.
  3. 3.
    We conduct a comparative performance analysis under realistic network conditions, quantifying trade-offs between latency, intelligibility, and computational load across deployment scenarios.
The remainder of this paper is organized as follows. Section II describes the acoustic signal modeling framework. Section III reviews the visual feature extraction and fusion strategies. Section IV presents the proposed AVSE architecture and deployment designs. Section V reports a comparative evaluation of deployment scenarios, and Section VI concludes with implications for future AVSE systems.
II Acoustic Signal Modeling
Central to the concept of AVSE is the nuanced modeling of acoustic signals. Here, each impulse response emanating from a spatial position of a target source to a microphone is represented as $h_{s}[k]$, where $k$ signifies a discrete-time index. This impulse response is bifurcated into two distinct components: the early part $h_{s}^{1}[k]$, encompassing the direct sound and early reflections, and the late part $h_{s}^{2}[k]$, which pertains to the subsequent reverberations. Consequently, the observed acoustic mixture at any given instance $k$ can be expressed in a mathematical formulation as:
$y[k]=\sum_{s=1}^{T}x_{s}[k]+d[k]$ (1)
$x_{s}[k]=x_{s}^{0}[k]\circledast h_{s}^{1}[k]$ (2)
$d[k]=\sum_{s=1}^{T}x_{s}^{0}[k]\circledast h_{s}^{2}[k]+\sum_{c=1}^{N}d_{c}[k]$ (3)
In this formulation, $x_{s}^{0}[k]$ represents the original speech signal emitted by the $s$-th target speaker. The term $x_{s}[k]$ denotes the clean speech signal from the $s$-th target speaker, inclusive of low-order reflections as captured by the microphone. The component $d_{c}[k]$ symbolizes the noise contribution from various sources as observed at the microphone.
III Visual Information Integration
The incorporation of visual data significantly augments the AVSE process. Studies, such as those mentioned in [25], have demonstrated the critical role visual elements play in enhancing speech recognition systems. Visual features not only provide high-level information about speech patterns but also furnish fine-grained articulatory details. The visual processing initiates with face detection and tracking algorithms [11], [17], [10], [4], [7], which are vital for delineating input dimensionality and discerning the count of speakers. This step significantly reduces the complexity of subsequent stages by focusing solely on the target faces.
III-A Deep Learning in AVSE
The actual fusion and processing of the audio and visual data in AVSE are achieved using an ensemble of deep neural network models. These models, by virtue of their design, can learn and synthesize representations of the acoustic and visual modalities at several levels of abstraction. Convolutional Neural Networks (CNNs) [6], [16], [18], [19], [24], with their inherent characteristics of sparse connectivity and translation invariance, are typically employed to process visual frames and extract pertinent features. They also find applications in acoustic signal processing, handling either the spectrogram or the raw waveform. The training of these networks, often conducted using stochastic gradient descent (SGD) [2], [1], RmsProp [12], and Adam [15], plays a critical role in parameter optimization and model effectiveness.
III-B Fusion Paradigms With Deep Learning
In the realm of AVSE, various strategies, including early and late fusion [23], [21], [22], are employed. Nonetheless, the predominant technique is intermediate fusion [20], where different unimodal representations amalgamate in a single hidden layer. This paper focuses on concatenation fusion, wherein audio and visual feature maps merge at corresponding spatio-temporal locations, allowing subsequent layers to decipher and integrate the information effectively. The fusion function, mathematically represented as $f:y_{l}[k],v_{l}[k^{\prime}]\rightarrow x_{l}[k]$, combines the audio feature map $y_{l}[k]$ and the video feature map $v_{l}[k^{\prime}]$ at times $k$ and $k^{\prime}$, respectively, producing an output map $x_{l}[k]$ at the $l$-th layer of the DNN. $x_{l}^{\text{cat}}[k]=f_{c}at(y_{l}[k],v_{l}[k^{\prime}])$ stacks the two feature maps at the same spatio-temporal locations $i,j,k$ across the feature channels $d$:
${x_{l}^{\text{cat}}[k]}_{i,j,2d}={y_{l}}_{i,j,d}[k]\quad\text{and}\quad{x_{l}^{\text{cat}}[k]}_{i,j,2d-1}={v_{l}}_{i,j,d}[k^{\prime}]$ (4)
where $x_{l}^{\text{cat}}[k]\in\mathbb{R}^{H\times W\times 2D}$.
Concatenation does not inherently define a correspondence but leaves it to subsequent layers to define, by learning suitable filters that weight the layers.
III-C Recurrent Neural Networks in AVSE
Given the inherently sequential nature of speech signals, Recurrent Neural Networks (RNNs) [5], Long Short-Term Memory networks (LSTMs) [5], [8], and Gated Recurrent Units (GRUs) [14] are particularly suitable for AVSE. These networks are adept at processing data where temporal dynamics are significant, making them ideal for speech signals where temporal coherence is paramount.
IV Proposed Solution
@@
Figure 1 presents a pragmatic framework for the deployment of the developed AVSE algorithm within the context of real-world applications. In order to enable the service in lowquality terminals (e.g., IoT devices), the remote server takes over all the processing, including the preprocessing of the raw data. By doing so, any device with a camera, microphone, and communication capabilities can use the service without any requirements on its computing power. This allows us to apply the service to some cheap IoT devices to achieve specific functions. For high-performance devices (e.g., smartphones), offloading all processing to the server can also reduce energy consumption and improve battery life.
@@ V Experimental Results
V-A Algorithm Processing Latency
@@
V-B Algorithm development
To mitigate the inherent latency in our system, we tried to refine our network’s proficiency in distilling potent feature representations from short segments of audio and visual data. To this end, we have calibrated the network training scheme to work with condensed input durations of merely $200$ ms. Leveraging the architecture depicted in LABEL:figure1, we constructed a series of networks, reducing their structures by paring down layers and neural nodes with the express aim of shortening processing delays. Table I lists the total parameters and processing latency of each of the three networks.
@@ V-B Algorithm development
VI CONCLUSION
In conclusion, this paper’s exploration of an Audio-Video-Based Speech Enhancement (AVSE) system underscores its potential to revolutionize speech recognition by integrating audio and visual data through advanced deep learning techniques. Among the myriad of practical applications, one of the most impactful is in the development of next-generation hearing aids. These advanced hearing aids, equipped with AVSE capabilities, could offer a transformative experience for individuals with hearing impairments, providing them with an enhanced ability to understand speech in noisy environments by leveraging both auditory and visual cues.
Looking to the future, several promising research directions could further elevate AVSE technology. Optimizing the balance between audio and visual modalities is crucial, as it can vary depending on environmental factors and the specific needs of the application. Investigating diverse DNN architectures can potentially unveil more effective ways of processing and merging audio-visual data. Moreover, experimenting with various fusion functions beyond simple concatenation could lead to more nuanced and context-aware integration of sensory inputs.
The broader implications of AVSE technology extend to a variety of real-world scenarios. Its application in military communications can drastically improve operational efficiency on the battlefield. In construction sites or any noisy industrial environment, AVSE can facilitate clearer internal communications among workers. The technology also has significant applications in videoconferencing, journalism, and sports, particularly in enhancing communication in noisy stadiums.
Each application presents unique challenges, suggesting a need for AVSE solutions that are customized to specific environments and user requirements. Future advancements in this field not only promise to push the boundaries of speech enhancement technologies but also offer substantial improvements in diverse sectors that rely on clear and effective communication. The integration of AVSE systems into hearing aids represents a particularly profound opportunity to enrich the lives of individuals with hearing impairments, marking a significant step forward in assistive technology.

### v2

@@ Abstract
This paper presents a real-time wireless 5G cloud-based audio-visual speech enhancement (AVSE) framework that integrates convolutional neural networks (CNNs) for acoustic feature extraction and OpenCV-based facial embedding for visual processing. The extracted multimodal features are fused through a long short-term memory (LSTM) network to reconstruct clean speech from noisy input. Experimental evaluations across multiple network configurations—Wi-Fi, 4G, 5G, Ethernet, and AWS cloud—demonstrate that only 5G and direct Ethernet links consistently satisfy the coherence condition ($t_{\text{comm}}\leq t_{\text{chunk}}$) for continuous streaming. By applying an 80× compression factor, the payload per 40 ms audio-video chunk is reduced from approximately 0.3 MB to 70 KB with minimal quality loss, enabling reliable transmission under bandwidth constraints. Results further reveal a trade-off between algorithmic latency and enhancement performance: smaller models and shorter input windows lower processing time but reduce speech clarity. The proposed architecture achieves end-to-end latency of approximately 5 s with strong noise reduction and speech intelligibility, demonstrating the feasibility of deploying multimodal speech enhancement in cloud-assisted, low-latency wireless environments.
@@
Audio-Visual Speech Enhancement, 5G Networks, Cloud Computing, Deep Learning, CNN-LSTM Fusion, Real-Time Processing, Wireless Communication, Multimodal Signal Processing
@@ I Introduction
In many real-world acoustic environments, speech signals are severely degraded by additive background noise, reverberation, and competing speakers, which substantially reduce intelligibility and perceived quality. Conventional audio-only speech enhancement techniques aim to recover a clean speech signal ($x[n]$) from a noisy observation ($y[n]=x[n]+d[n]$), where ($d[n]$) denotes noise or interference. However, when the signal-to-noise ratio (SNR) is low or overlapping talkers are present, these audio-only approaches often struggle to reliably recover intelligible speech.
Human listeners naturally exploit not only auditory cues but also visual information — such as lip movements and facial articulations — to compensate for challenging acoustic conditions. This perceptual phenomenon is well known (e.g., the McGurk effect) and underscores the value of multimodal input in speech perception. Visual speech cues remain unaffected by acoustic masking and therefore offer complementary information that is particularly valuable when the acoustic channel is corrupted.
Audio-visual speech enhancement (AVSE) builds on this insight by jointly leveraging both audio and visual modalities (for example, microphone audio plus camera capture of the speaker’s lips) to enhance or recover target speech. Deep-learning based AVSE systems typically extract acoustic features (e.g., log-mel spectrograms, STFT magnitudes) and visual features (e.g., lip region embeddings via CNNs), and then fuse them in a neural network to generate enhanced speech. Surveys (e.g., Michelsanti et al. 2021 [michelsanti2021overview]) have shown that multimodal AVSE systems often outperform audio-only baselines, especially in low-SNR or multi-speaker conditions.
Despite these algorithmic and modelling advances, deploying AVSE in real-world applications introduces a distinct set of system-level challenges. Wearable devices such as hearing aids or smart glasses must operate under tight constraints of latency, power consumption, and wireless connectivity. They also need to handle streaming audio and video data in real time. Offloading heavy deep‐learning inference to a remote cloud or edge server emerges as a feasible solution when the local device is resource‐constrained.
In this context, the advent of next-generation wireless networks (e.g., 5G NR) with high bandwidth and ultra-low latency offers a promising platform for real-time AVSE deployment. For example, recent demonstrations of 5G/IoT enabled AVSE for multimodal hearing aids illustrate the viability of high-data-rate, low-latency uplinks combined with cloud-based enhancement [gupta20235g]. By harnessing the 5G link and off-loading computation to the cloud, devices can remain light, efficient and power-conserving, while still benefiting from powerful multimodal models.
However, the full stack of implementation trade‐offs—including uplink/downlink latency, wireless reliability under mobility, compute placement (device vs edge vs cloud), streaming bandwidth for visual + audio, and real-time intelligibility/quality outcomes—remains under-explored in AVSE research. The present work addresses this gap by proposing and implementing a wireless 5G-based cloud architecture for AVSE. Our system captures lip-region video and noisy speech on a wearable front‐end, transmits the data via a 5G uplink to a remote cloud server, performs multimodal deep-learning AVSE processing, and streams back the enhanced speech to the device. We evaluate the system in terms of end-to-end latency, wireless throughput, model inference time, and speech intelligibility/quality under realistic noisy and multi-speaker conditions. In so doing, we extend the state of the art from model-centric AVSE research to a full-stack, deployment-oriented system suitable for mobile and wearable real-time use.
The remainder of this paper is structured as follows. Section II reviews related work on AVSE, wireless/5G off‐load implementations and deployment strategies. Section III presents the system design and implementation of our 5G cloud-based AVSE pipeline. Section IV reports experimental evaluation results and discusses findings and practical implications. Finally, Section V concludes the paper and outlines future research directions.
II Related Work
Research in audio-visual speech enhancement (AVSE) has progressed substantially over the past decade, particularly with the proliferation of deep learning and multimodal approaches. Early work in audio-visual speech separation, such as the seminal model by Ephrat et al. (“Looking to Listen at the Cocktail Party”), demonstrated that a joint audio‐visual network significantly outperforms audio‐only baselines in multi‐speaker and noisy conditions by leveraging lip‐motion cues [ephrat2018looking]. Subsequently, surveys such as Michelsanti et al. (2021) [michelsanti2021overview] provided a systematic overview of the field, highlighting key components such as acoustic feature extraction, visual feature encoding, fusion strategies (early vs intermediate vs late), and standard objective functions for AVSE/AVSS.
In more algorithm‐centric developments, attention has turned to how to fuse modalities effectively and design architectures that can operate in real‐time or near‐real‐time. For example, Xu and Hao [xu2021multi] proposed the attentional multi‐layer fusion convolution network (AMFFCN) to apply soft‐threshold attention at each layer of the decoder to select informative modalities. More recently, Gao and Grauman introduced the VisualVoice framework, which learns cross-modal embeddings (face appearance plus audio) for improved speech separation in unconstrained video scenes [gao2021visualvoice].
In terms of system implementation and deployment, particularly for wearable or real‐world applications, researchers have begun to tackle not just the algorithmic enhancement but also the architectural trade-offs of where the processing is done (device, edge, or cloud). Adeel et al. (2018) [adeel2018real] envisioned a 5G IoT enabled lip-reading driven hearing‐aid framework that off-loads AV data to the cloud, employing lightweight encryption and aiming for round‐trip latency targets on the order of a few milliseconds. On the dataset and baseline front, Michelsanti’s survey [michelsanti2021overview] underscores the importance of dataset design and evaluation metrics (e.g., STOI, PESQ, SI-SDR) when comparing AVSE systems.
Despite algorithmic advances, few works systematically compare deployment strategies (cloud vs edge vs device) under realistic network and computing constraints. To fill this gap, our work develops a comprehensive AVSE architecture (CNN + LSTM fusion) and performs a comparative analysis across deployment scenarios under Ethernet, Wi-Fi, 4G and 5G conditions. Although, prior work has established the foundational value of adding visual cues to speech enhancement; explored a wide variety of fusion architectures and benchmarks; and begun to address deployment challenges, The full spectrum of practical trade‐offs—end‐to‐end latency, network delay, compute location, and real‐world system performance—remains less explored, motivating the present implementation and experimental‐study contribution.
III Proposed Solution
@@
Figure 1 presents a pragmatic framework for the deployment of the developed AVSE algorithm within the context of real-world applications. In order to enable the service in low quality terminals (e.g., IoT devices), the remote server takes over all the processing, including the pre-processing of the raw data. By doing so, any device with a camera, microphone, and communication capabilities can use the service without any requirements on its computing power. This allows us to apply the service to some cheap IoT devices to achieve specific functions. For high-performance devices (e.g., smartphones), offloading all processing to the server can also reduce energy consumption and improve battery life.
@@ IV Experimental Results
IV-A Algorithm Processing Latency
@@
IV-B Algorithm development
To mitigate the inherent latency in our system, we tried to refine our network’s proficiency in distilling potent feature representations from short segments of audio and visual data. To this end, we have calibrated the network training scheme to work with condensed input durations of merely $200$ ms. Leveraging the architecture depicted in Figure 7, we constructed a series of networks, reducing their structures by paring down layers and neural nodes with the express aim of shortening processing delays. Table I lists the total parameters and processing latency of each of the three networks.
Figure 7: Audio-Visual Speech Enhancement AI model architecture.
@@ IV-B Algorithm development
V CONCLUSION
This study demonstrated a complete 5G cloud-based audio-visual speech enhancement (AVSE) system that integrates CNN-based acoustic processing and OpenCV facial feature extraction with an LSTM fusion network for temporal coherence. Experimental results showed that real-time performance depends critically on network latency and algorithmic processing time. Only 5G and wired Ethernet consistently met the coherence requirement ($t_{\text{comm}}\leq t_{\text{chunk}}$) (40 ms) for uncompressed 0.3 MB audio-video chunks, while compression at a factor of 80 reduced payloads to about 70 KB with negligible degradation, enabling smoother operation in constrained networks.
The findings highlight that network quality, chunk size, and model complexity jointly determine system responsiveness and enhancement quality. Although shorter input windows and smaller models reduce processing latency—from about 1.2 s to 0.35 s—they also degrade speech reconstruction, particularly under weak target signals. Achieving optimal performance therefore requires balancing temporal context and computational efficiency. Future work will focus on adaptive chunking, lightweight streaming architectures, and edge-assisted processing to further minimize delay while preserving enhancement fidelity in real-world wireless environments.
@@ Acknowledgements
This work was supported by the UK Engineering and Physical Sciences Research Council (EPSRC) Grant No. EP/XXXXXXX/X. The authors would like to gratefully acknowledge Haifeng Luo from the University of XXXX for his contribution in implementing XXXX.